OpenStudy (anonymous):

Use synthetic division to determine whether the number k is an upper or lower bound (as specified for the real zeros of the function f). k = 2; f(x) = 5x^4 + 4x^3 - 2x^2 + 2x + 4; Upper bound?

OpenStudy (anonymous):

@zepdrix could you help me with this one??

OpenStudy (anonymous):

@radar can you help!?

OpenStudy (anonymous):

@radar .................

OpenStudy (radar):

I am not familiar with "k" and bounds with 4th degree polynominals. Sorry. I will standby and hope someone will help you so I can learn what this is all about!!

OpenStudy (anonymous):

@radar thaankyouu by any chance, do you know how to do this one?? The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 30 + 0.006p3 and that its demand function is D(p) = 150 - 0.12p2, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (radar):

I don't know but I would set the functions to equal each other and convert to the equation .006p^3 + .12p^2-120=0 and use Wolfram. That would give you a real answer of 21.8586 for p plus two imaginary answers. As I mentioned anything above a quadratic is beyond my solving skills lol.

OpenStudy (anonymous):

haha thank you so much! @radar

OpenStudy (radar):

Sorry, but wish you luck with those.